Write the equation of variation.( inversely)
The number of copies produced by a copier varies directly as the amount of time the copier is operating. If the copier produces 150 copies in 4 minutes, how many copies can it produce in 1 hour?

Answer by ewatrrr(24785)

C = kt
0r
C/k = t
150/k = 4
150/4 = k
C/(150/4) = 1
C = (150/4) = 37.5